OpenStudy (anonymous):

factorise fully 8mp - 12

OpenStudy (anonymous):

8mp - 12 = 4(2mp - 3)

OpenStudy (anonymous):

should be 8mp-12=8mp-12 >8mp-8mp=+12-12 >8mp-8mp=0 ,then +12-12=0 i think its right.

OpenStudy (anonymous):

i think you've confused factorising with solving equations. when we factorise we want to put something into brackets, like 45 = 9(5) if theres no equals sign in what you're given, we cant really solve it like you tried to do (i think thats what you're doing) we are looking for common factors; what is a factor of 8mp AND -12 ? well both can be divided by 4 so what we do is divide them by four and put it outside the bracket: 8mp - 12 = 4(2mp - 3)

OpenStudy (anonymous):

do you see what i mean?

OpenStudy (anonymous):

ah,i see.